OpenStudy (anonymous):

I need help with 4 questions!

OpenStudy (anonymous):

those are ?

OpenStudy (anonymous):

I NEED HELP WITH 3 QUESTIONS ASAP

OpenStudy (anonymous):

Solve for h. P = mgh A. h = P – mg B. H=pg/m C.h=p/mg D. h = P + mg

OpenStudy (anonymous):

I have no clue `-` what grade are you in?

OpenStudy (anonymous):

7th

OpenStudy (anonymous):

e.e I'm in k12 I'm in 8th and i don't even know

OpenStudy (anonymous):

I'm i K12 too!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

good luck i bet i didn't past this

OpenStudy (anonymous):

haha I'll open up a new question thanks for trying anyways :)

OpenStudy (anonymous):

no probelm and tr to get owl bucks those will help you out cya later